OpenStudy (anonymous):

simplify by factoring assume that all expressions under the radical represent non-negative numbers. ^4sqrt32x^12y^14

OpenStudy (anonymous):

\[\sqrt[4]{32x^{12}y^{14}}\] \[=\sqrt[4]{2^4\times 2\times (x^3)^4\times( y^3)^4\times y^2}\] \[=2x^3y^3\sqrt[4]{2y^2}\]